I have been working on this problem all day. can I get some help with this problem:
-4(x+2)+6=2(x+3)+4

Answer by huntingmaster325(15)

-4(x+2)+6=2(x+3)+4 for cant do nothing in the ( )
but multiply but multiply what is out side the ( ) to what in the ( )
-4x+-8+6=2x+6+4 not put what incommon together/
-4x-2=2x+10 add -4x to each side and subtract 10 from each side.
-12=6x divide both side by 6
-2=x
answer is -2=x